Operations Report – December 27th

Crew 218 Operations Report 27-Dec-19

SOL: 05
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 13
From what time last night: 1930
To what time this morning: 0830
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before the generator is run at night: 76 %
Diesel Reading – 83%
Station Propane Reading – 78%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 43
Water Meter (units): 0147061.8
Water (static tank) (gal): 410
Static to Loft Pump used – Yes
Water in Green Hab (gal): 110
Water in Science Dome (gal): 0
Toilet tank emptied: No

Deimos rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Not used
Hours: 119.2
Beginning charge: (Before EVA):
Ending charge: (On return from EVA, before recharging):
Currently charging: Yes
Opportunity rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Not used
Hours: 124.4
Beginning charge:
Ending charge: (On return from EVA, before charging):
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A
HabCar used and why, where? No
CrewCar used and why, where? No
General notes and comments: N/A
Summary of the internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: The pipe from the static tank through the loft tank frozen. The crew rationed the water during the morning until 3:00 PM that the water in the pipes melted. The water system from the Static tank to the Loft tank is working as usual.
Summary of GreenHab operations: routine watering. Moved hanging plants: snow peas and buttercrunch away from heater vent. They are now hanging over the herbs. Recommend transplanting herbs to pots, and transplanting larger plants (tomatoes, peppers, etc.) to Plot E.

Summary of Science Dome operations: Nominal
Summary of RAM operations: Spare fire alarm in the RAM storage closet went off at 12.20 PM and 12.40 PM. HSO removed the battery.
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns, and requests to Mission Support: NA

Luz Ma

ENG

Crew 218